OpenStudy (anonymous):

I need help the components of vector a is ax = 2.2 and ay = -6.9, and the components of vector b are given bx = -6.1 and By = -2.2. What is the magnitude of the vector b - a?

OpenStudy (anonymous):

Use this formula:- |b-a|=sqrt((bx-ax)^2+(by-ay)^2)) You will get 9.538

OpenStudy (anonymous):

Hi Akki, I'm coming up with 9.9

OpenStudy (anonymous):

B = -6.1 + -2.2 - A = 2.2 + -6.9 ---------------- -3.9i + -9.1j sqrt ((-3.9)^2 + (-9.1)^2) = 9.9 Please assist, thanks

OpenStudy (anonymous):

No Juan, I'm trying to figure out how Akki got 9.5. Thanks for any help you can provide

OpenStudy (e.mccormick):

Here is an overview of vector components: http://www.physicsclassroom.com/class/vectors/Lesson-1/Vector-Components